OpenStudy (anonymous):

3x=sqr of 9x+2 over 2 been stuck on this one for a while

OpenStudy (matt101):

\[3x=\sqrt{(9x+2)/2}\] \[9x^2=(9x+2)/2 \] \[18x^2=9x+2 \] \[18x^2−9x−2=0 \] \[(3x−2)(6x+1)=0 \] \[x=2/3,−1/6\] However, remember that we started with a radical, and a radical can only produce a positive number. If we use x=-1/6, we see the radical must equal -1/2 (=3x), which is not possible. Therefore, the only real solution is 2/3. The reason why there's a second solution is because as part of solving the equation, I squared both sides, introducing a second possible solution.

OpenStudy (anonymous):

:) oh thank you!